How many joules are absorbed as the temperature of 15.0 grams of aluminum is heated from 20.0oC to 60.0oC?



1
Expert's answer
2021-10-28T02:57:54-0400

"Q = mcΔT \\
ΔT = 60.0-20.0 = 40.0 \\
m = 15.0 \;g \\
c_{Al} = 0.900 \;J/g \cdot C \\
Q = 15.0 \times 0.900 \times 40.0 = 540 \;J"